OpenStudy (anonymous):

How many different strings can be made from the letters in ORONO, using all or some letters?

OpenStudy (anonymous):

I need to know how to do this using combinations and stuff

OpenStudy (anonymous):

Or maybe it's permutations.

OpenStudy (anonymous):

rather a pain since you have to do it for one, two, three, four, and five

OpenStudy (anonymous):

for 1 letter there are only 3

OpenStudy (anonymous):

for 2 letters it looks like there are 7 different ones OR RO ON NO RN NR OO

OpenStudy (anonymous):

five letter strings are easiest, since there are \[\frac{5!}{3!}=5\times 4=20\] strings

OpenStudy (anonymous):

i think this actually requires a bit of work for 3 and 4, although the answer is 13 and 20 respectively, so the total is 3+ 7 + 13 + 20 + 20 = 63 a tree diagram would be the best way probably

OpenStudy (anonymous):

So, a tree diagram is the best way?